Run Planning Meeting Notes, June 8 - June 15, 2022

From GlueXWiki
Jump to: navigation, search

Meeting are via ZoomGov: connect (Meeting ID: 160 199 0060 / Passcode 8394)

MCC whiteboard

Hall D counting room white board

Tuesday June 14

  1. Last 24h:
    • 7am-7:40am: CToF HV scan
    • 7:40am-8:15am: "production" on Amo 3E-4 RL, 100 nA.
    • 8:15am-10:30am: Beam taken away for optic study in the machine.
    • 11am-7:15pm: Diamond alignment.
    • 7:15pm-7:30pm: Ion Chamber calibration on diamond.
    • 7:40pm-00:10am Pb absorber study.
    • 00:10am-00:40am Hall D preparing to go back to "production".
    • 00:40am-5:20pm start overnight production on Amo 3E-4 RL, 60 nA. (Current reduced from 100nA because of low livetime).
    • Beam down. Opportunistic access to install steel shielding for CToF.
    • 6:40am-7:40am: "Production runs" to see effect of steel shield and check livetime problem.
    • 7:40am beam down for maintenance, for at least 8h.
    • Requested tagger in control access for investigating TAGH alarm issue and NMR probe.
  2. Reports
  3. Issues
  4. Plans
    • Access to install steel shielding plate downstream of CTOF (already done).
    • Access for investigating TAGH alarm issue and NMR probe.
    • When beam comes back (swing shift?)
      • Harp scans.
      • Flux optimization/Act .Col. calibration
      • NPP production run overnight (even if previous item failed). Diamond, 45/135, 30nA or more possible (it will be up to NPP persons to check and tell if a higher current may be used).
  5. Commissioning steps temporarily postponed:
    • optimize photon beam transmission
    • PS and TAGH HV scans

Monday June 13

  1. Last 24h:
    • 7:30am-9:00am Microscope HV scans. Richard could not finish it before beam went down (see next item)
    • Because of Ops cannot deliver high current beam to Hall C and MPS trip issues, it was decided to have a 3h access in the injector in the morning. It starts around 9am.
    • The Monday injector access last well beyond 3h. The maintenance day of Tuesday disappears from the schedule.
    • 11am-13:30pm: control access in the tagger (Nick, to look at a TAGH board possibly causing alarms. Tim to look at the tagger magnet NMR that is no locking.).
    • 5pm. Hall D in controlled access (rapid). Malte went in to change FCal bases.
    • Beam came back at 23:15. Its convergence looks fine. It is too late for doing L1 trigger studies and lead absorber studies. The postponed them. Instead, we will take data overnight with new configuration file excluding CToF from the trigger, which permitted us to run at 100nA with 99.7% livetime and 12kHz DAQ rate, and do the microscope and CToF HV scans early in the morning.
    • 5:30am Richard starts finalizing the microscope HV.
    • 7:00am Beni start CToF HV scan (the one done Sunday night was not usable).
  2. Reports
    • MCC PD-RC meeting:
      • Despite the 14h beam down time for diagnostics, Ops and CASA still don't know the origin of the problem preventing Hall C beam delivery and causing MPS trip.
      • Hall D RC pointed out that with expert waiting for the beam to come all day, we could not commission during the night which means we had to take so not that useful data on amorphous radiator.
      • Despite not having resolved the beam problem PD decided to not repeat today the beam down diagnostic and let Hall B and D run so that:
        • Hall B can be ready to move to 5-pass tomorrow (done with their 1-pass program, will stay at 5-pass the rest of the run).
        • Hall D can align their diamond.
      • I expect beam will be down tomorrow for diagnostic (Hall C not able to run is a major issue) and pass change.
      • Hall D RC did not bring up the topic of using the B slit, because PD and ops have already enough on their mind, and bleedthrough is not an issue at the moment.
    • TOF HV adjustements.
    • Initial FWMPC timing alignment
    • Microscope HV scan (second and final part)
    • Radiation level studies
  3. Issues
  4. Plans
    • Early morning: CToF HV scan.
    • Microscope HV scan (Done).
    • Beam stability was good: diamond alignment today, starting just after the meeting.
    • Ion Chamber calibration on diamond if alignement is done.
    • If not too late in the afternoon: L1 trigger study if diamond alignment.
    • Pb absorber study.
    • Overnight production on amo, 3E-4 RL, 100nA.

Sunday June 12

  1. Last 24h:
    • 7:20am: Access to switch from Pb to empty target. Then no beam up to 8:20 due to accelerator having a power supply issue.
    • 8:20am-11:20am: DAQ problem (running at very low rate). Sergey investigated and fixed the issue.
    noon-2:40pm: Beam down (big thunderstorm over Newportnews).
    • Took about 2.5h of data on "empty target" (less than the planned 4h because of above problems).
    • 6:15pm-9:15pm: Beam down for Hall D due to some Hall C beam delivery issue that needed injector diagnostics. -> it became too late for L1 study. We postponed it to tomorrow morning.
    • 6:20pm Access to switch back to Pb target.
    • 30min run by Lubomir, with trigger without bit-6 to diagnose FDC occupancy problem.
    • Short raw mode run by Beni to diagnose CToF problem.
    • 10pm-midnight: CToF HV scans.
    • "Production" on Pb target overnight (60nA).
    • 5:30am-7:30am Beam down due to vacuum issue.
    • 7:30: Starting microscope HV scan.
  2. Reports
  3. Issues
  4. Plans
    • microscope HV scan, if the beam comes back before 9am
    • Beam will be down for 3h-4h (until past noon?) for injector diagnostic to fix beam trip problems.
    • L1 trigger work (A. Somov ~2h)
    • Adjust Pb absorber upstream of muon detector, first part (Ilya, ~5h). Will be completed tomorrow evening.
    • Finalize Microscope HV scan (~1h, Richard Jones).
    • Production overnight until beam is taken away for maintenance Tuesday morning.
  5. Commissioning steps temporarily postponed:
    • CToF HV scan
    • optimize photon beam transmission
    • Ion Chamber calibration on diamond
    • PS and TAGH HV scans

Saturday June 11

  1. Last 24h:
    • 8:40am: Finished ToF, (CToF) and CDC HV scans.
    • 8:45-12:50: Beam down due to injector issue.
    • Beam convergence after beam down time looks good
    • Did 15 min raw mode run 100 nA, Amo 3E-4 RL and using the configuration file Sergei prepared in the morning (part of detector checkout).
    • Attempt to optimize photon beam transmission. Aborted because beam position could not ne locked on the profiler.
    • L1 trigger set-up.
    • 20min raw mode run (Asked for by Beni to understand the CTOF rates / threshold settings)
    • Overnight: production on Amo 3E-4RL and Pb target, with improved trigger. (high rates in the CTOF => low lifetime ~50%, with beam current limited to 60 nA).
    • 5:30am-8:20am beam down due to power supply issue.
    • While te beam is down, access to switch from Pb target to "empty" target.
  2. Reports
  3. Issues
  4. Plans
    • 4h data taking on empty target. Then, short access (Rory) to put back the Pb target.
    • 20 min run on production mode, but without trigger bit 6. (Lubomir)
    • L1 trigger work (A. Somov)
    • Short raw mode run (Beni)
    • CToF HV scan (Beni)
    • Production on Pb target overnight.
  5. Commissioning steps temporarily postponed:
    • microscope HV scan (Monday?)
    • CToF HV scan
    • Ion Chamber calibration on diamond
    • PS and TAGH HV scans

Friday June 10

  1. Last 24h:
    • Beam was down for about 10h (~9am-7pm) due to RF zone problem and then electrical panel breaker work.
    • Short Hall access by Nick Sandoval.
    • When beam came back, Harp scans showed that beam convergence was good enough for the next runplan steps:
      • A. Somov did some preliminary L1 trigger studies. Was not able initially to run at the desired 100nA. Could not run in raw mode.
      • Detector calibration. Did the 1h production mode run (at 60nA instead of 100nA), but the 15min raw mode run was postponed (see above).
      • Beni did his ToF and CToF HV scans. Was able to do it at 100nA. CToF HV will need to be redone.
      • Beni/Night/morning shift crews did CDC HV scan, following instructions written by Naomi.
  2. Reports
  3. Plans
    • Check beam convergence.
    • Do a 15min raw mode run with the configuration file prepared by Sergei.
    • Finalize GlueX+CToF trigger. Fix problem with raw mode running. Preliminary look at CPP/NPP trigger
    • If Sacha has time: do the PS/TAGH HV scan.
    • Take 1h run with the new trigger setup. (If it is late in the evening, make this overnight runs)
    • If it is not too late in the evening: Optimize beam transmission and Act. col. calibration.
    • Overnight runs with the new trigger setup.

Thursday June 9

  1. Last 24h:
    • Ion Chamber calibration (had to be redone this morning as OPS found the one done yesterday suspicious).
    • Radiation levels checks done (levels, according to our CARMS, are very good).
    • Radiation levels scans with current (50nA, 75nA, 100nA)and RL (0, 1E-4, 3E-4) done. Need to be analyzed.
    • Microscope HV scan postponed to Monday. (Microscope can still be used.)
    • Most of the time was spend on tuning beam and doing harp scans (5C11, 5C11b, ours). Beam was not good, with bad convergence/divergence and showing significant background. May have been improved during the night (needs to be checked).
    • Act. Col. operational.
  2. Reports
  3. Plans
    • Check photon beam quality (Hovanes, already started around 8am.)
    • Do L1(GlueX)/DAQ/detector checkout and take a bit of data?
    • Align diamond during the day? (Might not be the best time: Hall C is expecting to finish its access around 9:30. Then MCC will resume high current tune to Hall C).
    • Accelerator may be down for ~2h today to bring back a RF cavity. No firm plan yet.
    • Do ToF/CToF/CDC HV scans in the evening?

Wednesday June 8

  1. Last 24h:
    • Ops is tuning the beam and asked Hall D to be ready by 11am-noon for physics beam. We restored shift starting at noon. But beam is coming only just now (7am-8am, Thursday 9th). Doing Ion Chamber calibration
      • Lost a few hours because of repair work done on a fan dropped a zone (miscommunication with Facility).
      • Struggled to tune Hall C and this put Hall D tune on backburner. Wasted 8-10h (whole night) trying to tune Hall C with two magnets off, which is not possible. DC power group is e-assessing their HCO procedure.
      • One of the quadrupole (MQP5C09) on Hall D line tripped on overtemperature. Probably due to clogged cooling line. Needed to investigate so brought down North Linac and ramped down our tagger magnet. This quad. is now off and will (tentatively) be repaired at next maintenance day. CASA devised a tune for us that does not use this quads.
    • Yesterday MCC safety minute: The PD saw a water mocassin eating a bullfrog in his pound => watch for snakes.
  2. Reports
    • A. Somov replaced two boards during late afternoon (rapid access).
    • If you need access, put your name on Hall D white board (access list) and let RC know.
  3. Issues
    • Maybe problem with 2-spot beam (two beams differing in energy by 10^-3). Not clear if this is present with CW beam. If it is, fixinf this may delay the program: fixing this is of the highest priority for Ops.
    • While walking to MCC yesterday (Wed. June 8th) afternoon, I passed a snake that was basking in the sidewalk gutter => watch for snakes.
  4. Plans
    • Ops still have some work to do: Establish high current in Hall C. Start Hall B (tomorrow?).
    • Assess beam quality, optimize beam transmission, checkout/calibrate systems.